# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v2

@@ 7.2 Stellarator physics
Stellarators have a unique external control over the plasma through the coils, which (Boozer:2004) showed could efficiently produce approximately fifty independent magnetic field distributions, an order of magnitude more than can be produced in axisymmetry. By their nature, stellarators provide steady-state plasma confinement, low recirculating power, as well as robustness against disruption and the loss of positional control within the plasma chamber.
Adequate stellarator confinement requires optimization of the external field. In 1980 stellarators were known to have many attractive properties for fusion power plant, but it was thought that this was precluded by the extremely large neoclassical transport caused by their toroidal asymmetry. (Boozer:1980) gave a new formulation of the drift equations and a new coordinate system that allowed a relatively fast optimization to reduce the neoclassical transport. (Nurenberg:1988) used this formalism to show that such magnetic configurations exist. The result of modern optimization for a fusion power plant is illustrated by the Infinity Two fusion pilot plant design (Hegna:2025).
@@
(Boozer:2024:Stellarators) gave a conceptual design for stellarators with enhanced tritium confinement and edge radiation control. This design used special features of the stellarator to greatly enhance the tritium burn-up fraction, a non-resonant divertor to allow impurity radiation at the edge to spread the thermal power exhaust broadly over the wall, and flattop density and temperature profiles to enhance the fusion power production and reduce the impurity confinement. The low tritium burn-up fraction is a major issue in obtaining tritium self-sufficiency in toroidal fusion power plants. It should be noted that if a fission source of tritium could be identified that allowed a large fusion to fission power ratio and with a cost of only millions of dollars per kilogram, fusion power plants could be greatly cheapened and simplified by reducing or eliminating their tritium-breeding blankets. A large fusion/fission power ratio would allow the fission tritium producers to be located at secure sites, such as nuclear weapons facilities. Poor confinement in the central part of the stellarator would flatten all profiles, including those of impurities, and as discussed in Section 6.4 the magnetic field chaos produced at non-zero plasma pressure by what is called electrostatic microturbulence could expel impurities.
7.3 Program priorities
The Merriam-Webster dictionary gives two definitions of innovation: (1) a new idea, method, or device and (2) the introduction of something new. By either definition, innovations cannot be explicitly predicted. Nevertheless, they are a central element in the achievement of practical fusion energy on the shortest possible timescale as well as advancing other important applications of plasma physics, such as space weather. Admiral Rickover led the development of the nuclear submarine, the Nautilus, which began operations approximately fifteen years after the splitting of the uranium nucleus in Berlin. His insights can provide guidance for the rapid achievement of fusion: “All new ideas begin in a non-conforming mind that questions some tenet of the conventional wisdom.” “We shrivel creativity by endless frustrations." “Good ideas are not adopted automatically. They must be driven into practice with courageous impatience. Once implemented they can be easily overturned or subverted through apathy or lack of follow-up, so a continuous effort is required." Rickover did not seek to eliminate the persons who were constantly developing new ideas from the program because a precise time line for developing new ideas could not be given.
The United States Department of Energy (US-DoE:2025) released the Fusion Science and Technology Roadmap, which prominently featured “Innovate and advance the science and engineering of fusion with well-defined milestones and metrics" on page 7. The inconsistency of innovations “with well-defined milestones and metrics" could be viewed as a wording issue but, whether intended or not, seemingly defines policy.
The reviewers of my April 2025 proposal on magnetic reconnection recognized the “various excellent contributions to reconnection research" that I had made but felt that continuing innovations without well-defined milestones, by the instructions they had received, only justified a low funding priority. This position was accepted by the Department of Energy. Recent papers on reconnection (Boozer:2025:Constraints) and (Boozer:2025:Electron) had no external support.
My proposal seeking continued support for “Non-axisymmetric magnetic fields in tokamaks” was submitted on 4 January 2024 and declined on 1 December 2025 for not having clear milestones despite very positive reviews. During the twenty-three months the renewal was under review, I had six publications relevant to the proposal. Four have been discussed in this article: (Boozer:2024:Electric), (Boozer:2025:Constraints), (Boozer:2025:Efficient), and (Boozer:2025:Magnetic). Two others are this article, and (Boozer:2024:Use), which provided a method for studying the unexplored design freedom of tokamak and stellarator coils. Eight additional publications during this period were on stellarators and the article (Boozer:2025:Electron). Could any of these contributions been foreseen for the three to five years needed to obtain grant approval when “well-defined milestones and metrics” are required?
My individual grants have expired, and only one renewal proposal remains. It is on stellarators and will also fail if past innovations are taken as having no importance for future funding. The issue is not just about me. Many group leaders as well as researchers feel too entangled in the spider web of milestones and metrics to do what they know is of greatest importance or to speak out. My situation makes it incumbent for me to try change a very destructive Department of Energy policy. If I am forced to retire, my taxable income will increase by fifty percent. This and my continuing, impactful, and well documented contributions allow me to publicly seek a change. My longterm contributions and their programmatic impact are documented in part at “Boozer-Google Scholar” with many publications single author.
Does anyone actually think that the achievement of fusion energy is made faster and less costly by removing grant support from persons whose research is focused on innovations that satisfy the standard meaning of being too novel to predict?
@@ Appendix A Separation of neighboring lines
The behavior of the field lines that are separated by a distance $\rho$ from an arbitrary line $\vec{x}_{0}(\ell)$ has a general form, (Boozer:2012), as $\rho\rightarrow 0$. The trajectories of these infinitesimally separated lines are given in Courant and Snyder intrinsic coordinates by a simple Hamiltonian. This coordinate system, which is defined in Appendix B of (Courant:1958), is $\vec{x}(\rho,\alpha,\ell)=\rho\cos\alpha\hat{\kappa}_{0}+\rho\sin\alpha\hat{\tau}_{0}+\vec{x}_{0}(\ell)$, where the $\hat{b}_{0}\equiv\vec{B}_{0}(\ell)/B_{0}$ along the line, $d\hat{b}_{0}/d\ell=\kappa_{0}\hat{\kappa}_{0}$ is the curvature of the line, and $d\hat{\kappa}_{0}/d\ell=-(\kappa_{0}\hat{b}_{0}+\tau_{0}\hat{\tau}_{0})$ gives the torsion $\tau_{0}$. The torsion measures the extent to which the line fails to lie in a plane. Letting $\tilde{\psi}\equiv\pi B_{0}(\ell)\rho^{2}$, the neighboring trajectories are given by the Hamiltonian
@@ A.1 Maximum field-line separation
is sometimes used to defined chaos, but $\lambda_{L}$ can be zero with $\sigma_{max}(\ell)$ reaching arbitrarily large values. For reconnection the maximum $\sigma_{max}(\ell)$ is what is relevant not $\lambda_{L}$, which may be zero.
@@ A.2 Dependence on only two coordinates
with $K_{\omega}$ and $k_{q}$ constants. The correctness of the evolution equations $d\alpha_{q}/d\ell=\partial\tilde{\mathcal{H}}/\partial\tilde{\psi}$ and $d\tilde{\psi}/d\ell=-\partial\tilde{\mathcal{H}}/\partial\alpha_{q}$ are easily checked. The generating function $G=(\alpha-\ell/R_{q})\tilde{\psi}$ transforms the canonical coordinate $\alpha$ to $\alpha_{q}=\partial G/\partial\tilde{\psi}$, transforms the Hamiltonian to $\tilde{\mathcal{H}}=\tilde{H}+\partial G/\partial\ell$ or $\tilde{\mathcal{H}}=\tilde{H}-\tilde{\psi}/R_{q}$, and leaves the canonical momentum $\tilde{\psi}$ unchanged.

### v3

@@ 6.3 Stellarator physics
Stellarators have a unique external control over the plasma through the coils, which (Boozer, 2004) showed could efficiently produce approximately fifty independent magnetic field distributions, an order of magnitude more than can be produced in axisymmetry. By their nature, stellarators provide steady-state plasma confinement, low recirculating power, as well as robustness against disruption and the loss of positional control within the plasma chamber.
Adequate stellarator confinement requires optimization of the external field. In 1980 stellarators were known to have many attractive properties for fusion power plant, but it was thought that this was precluded by the extremely large neoclassical transport caused by their toroidal asymmetry. (Boozer, 1980) gave a new formulation of the drift equations and a new coordinate system that allowed a relatively fast optimization to reduce the neoclassical transport. (Nürenberg and Zille, 1988) used this formalism to show that such magnetic configurations exist. The result of modern optimization for a fusion power plant is illustrated by the Infinity Two fusion pilot plant design (Hegna et al., 2025).
@@
(Boozer, 2024b) gave a conceptual design for stellarators with enhanced tritium confinement and edge radiation control. This design used special features of the stellarator to greatly enhance the tritium burn-up fraction, a non-resonant divertor to allow impurity radiation at the edge to spread the thermal power exhaust broadly over the wall, and flattop density and temperature profiles to enhance the fusion power production and reduce the impurity confinement. The low tritium burn-up fraction is a major issue in obtaining tritium self-sufficiency in toroidal fusion power plants. It should be noted that if a fission source of tritium could be identified that allowed a large fusion to fission power ratio and with a cost of only millions of dollars per kilogram, fusion power plants could be greatly cheapened and simplified by reducing or eliminating their tritium-breeding blankets. A large fusion/fission power ratio would allow the fission tritium producers to be located at secure sites, such as nuclear weapons facilities. Poor confinement in the central part of the stellarator would flatten all profiles, including those of impurities, and as discussed in Section 6.1.4 the magnetic field chaos produced at non-zero plasma pressure by what is called electrostatic microturbulence could expel impurities.
7 Discussion
The incentive for a physicist to learn an area of mathematics is insights on and methods for solving problems of practical importance. Examples show that the lack of familiarity with the mathematical concepts of magnetic field line chaos, cantori and turnstiles block developments in toroidal plasma physics. This was the motivation for this review, which is an expanded version of an invited talk at the 29th Workshop on MHD Stability Control in July 2025.
The importance of field line chaos to the breaking of magnetic field line connections may seem obvious, but the concept receives little notice at international conferences on reconnection. The work of (Borgogno et al., 2011) on magnetic reconnection is an exception. They considered effects of cantori and turnstiles without using the names—concepts that could have made the generality of their results more apparent. The ease and rapidity of magnetic reconnection when the magnetic field is chaotic implies what are called electrostatic instabilities can produce non-ambipolar transport through an electron viscous force.
The magnetic energy that is released by the breaking of connections in chaotic magnetic fields requires an understanding of the damping of Alfvén waves in chaotic magnetic fields, which are the immediate recipients of that energy. Phase-mixing arguments correctly predict this damping is rapid, but the fraction of the power that is dissipated by resistivity, which is primarily a transfer to the electrons, versus that dissipated by viscosity, which is primarily a transfer to the ions, is not understood.
The primary definition of chaos in the Merriam-Webster dictionary is “a state of utter confusion,” but a definition using technical terms is also given: “the inherent unpredictability in the behavior of a complex natural system.” Both make the formation of highly collimated flux tubes implausible when the magnetic field is chaotic. Nevertheless, this collimation is the essential element in the explanation of the extreme localization in space and time of runaway electrons strikes on the chamber walls. This localization presents the greatest danger to tokamak devices following a disruption. The successful avoidance of this destructive collimation by (Reux et al., 2021) and (Paz-Soldan et al., 2021) using a rapidly-developing instability is counterintuitive without an understanding of cantori and turnstiles.
Non-resonant divertors offer the potential for far more control of the plasma edge than do divertors based on a separatrix. Control is needed for the out-coming particles to follow well-collimated flux tubes into the highly localized ducts of pumps while defining a sufficiently thick region of poor confinement for high-Z impurities to spread the out-coming power over the walls and and to block in-coming neutrals from charge exchanging within the high energy plasma. Charge exchange would produce high-energy neutrals that erode the walls, Boozer (2024b). Compared to their potential for solving a major issue in the development of fusion power plants, little research is being done on non-resonant divertors, which requires practical methods for controlling chaos, cantori, and turnstiles.
When there is a pressure gradient along magnetic field lines, $\vec{B}\cdot\vec{\nabla}p\neq 0$, in addition to the pressure gradient across the field, $\vec{B}\times\vec{\nabla}p$, quasi-neutrality becomes far more subtle than ambipolarity. Ambipolarity is enforced by a constant electric potential on each magnetic surface. Although generally ignored in simulations, this subtlety always arises in divertors and during disruptions. Subtleties exist even without magnetic field line chaos, but with chaos the $\vec{E}\times\vec{B}$ drifts required by quasi-neutrality give transport at a Bohm-like rate. Practical methods for including these effects in simulations require development, and the realism of effects seen in simulations, which have far less spatial and temporal resolution than is in principle required, remains to be determined.
The literature on the applications of chaos, cantori, and turnstiles in toroidal plasma physics is concentrated among only a few authors, which implies much could be gained from a broader understanding of these concepts.
@@ Appendix A Separation of neighboring lines
The behavior of the field lines that are separated by a distance $\rho$ from an arbitrary line $\vec{x}_{0}(\ell)$ has a general form, (Boozer, 2012), as $\rho\rightarrow 0$. The trajectories of these infinitesimally separated lines are given in Courant and Snyder intrinsic coordinates by a simple Hamiltonian. This coordinate system, which is defined in Appendix B of (Courant and Snyder, 1958), is $\vec{x}(\rho,\alpha,\ell)=\rho\cos\alpha\hat{\kappa}_{0}+\rho\sin\alpha\hat{\tau}_{0}+\vec{x}_{0}(\ell)$, where the $\hat{b}_{0}\equiv\vec{B}_{0}(\ell)/B_{0}$ along the line, $d\hat{b}_{0}/d\ell=\kappa_{0}\hat{\kappa}_{0}$ is the curvature of the line, and $d\hat{\kappa}_{0}/d\ell=-(\kappa_{0}\hat{b}_{0}+\tau_{0}\hat{\tau}_{0})$ gives the torsion $\tau_{0}$. The torsion measures the extent to which the line fails to lie in a plane. Letting $\tilde{\psi}\equiv\pi B_{0}(\ell)\rho^{2}$, the neighboring trajectories are given by the Hamiltonian
@@ A.1 Maximum field-line separation
is sometimes used to define chaos, but $\lambda_{L}$ can be zero with $\sigma_{max}(\ell)$ reaching arbitrarily large values. For reconnection the maximum $\sigma_{max}(\ell)$ is what is relevant not $\lambda_{L}$, which may be zero.
@@ A.2 Dependence on only two coordinates
with $K_{\omega}$ and $k_{q}$ constants. The correctness of the evolution equations $d\alpha_{q}/d\ell=\partial\tilde{\mathcal{H}}/\partial\tilde{\psi}$ and $d\tilde{\psi}/d\ell=-\partial\tilde{\mathcal{H}}/\partial\alpha_{q}$ is easily checked. The generating function $G=(\alpha-\ell/R_{q})\tilde{\psi}$ transforms the canonical coordinate $\alpha$ to $\alpha_{q}=\partial G/\partial\tilde{\psi}$, transforms the Hamiltonian to $\tilde{\mathcal{H}}=\tilde{H}+\partial G/\partial\ell$ or $\tilde{\mathcal{H}}=\tilde{H}-\tilde{\psi}/R_{q}$, and leaves the canonical momentum $\tilde{\psi}$ unchanged.